# Adapting Short-Term Transformers for Action Detection in Untrimmed Videos

### v1

@@ Abstract
Vision transformer (ViT) has shown high potential in video recognition, owing to its flexible design, adaptable self-attention mechanisms, and the efficacy of masked pre-training. Yet, it still remains unclear how to adapt these pre-trained short-term ViTs for temporal action detection (TAD) in untrimmed videos. The existing works treat them as off-the-shelf feature extractors for each short trimmed snippet without capturing the fine-grained relation among different snippets in a broader temporal context. To mitigate this issue, this paper focuses on designing a new mechanism for adapting these pre-trained ViT models as a unified long-form video transformer to fully unleash its modeling power in capturing inter-snippet relation, while still keeping low computation overhead and memory consumption for efficient TAD. To this end, we design effective cross-snippet propagation modules to gradually exchange short-term video information among different snippets from two levels. For inner-backbone information propagation, we introduce a cross-snippet propagation strategy to enable multi-snippet temporal feature interaction inside the backbone. For post-backbone information propagation, we propose temporal transformer layers for further clip-level modeling. With the plain ViT-B pre-trained with VideoMAE, our end-to-end temporal action detector (ViT-TAD) yields a very competitive performance to previous temporal action detectors, riching up to 69.0 average mAP on THUMOS14, 37.12 average mAP on ActivityNet-1.3 and 17.20 average mAP on FineAction.
@@ 2 Related Work
Figure 2: Overview of ViT-TAD. We suppose the ViT-based backbone has $n$ blocks and divide them into several subsets. Each subset has $i$ blocks. We divide a video clip into several snippets and send each into the backbone for feature extraction. We perform temporal feature interaction among all snippets through the inner-backbone information propagation strategy. We further conduct clip-level modeling to refine clip-level features through the post-backbone information propagation strategy. *** means the last layer is initialized as zero.
@@
[…]
@@ 3.1 Inner-backbone Information Propagation
We design a cross-snippet propagation module for inner-backbone information exchange. Unlike hierarchical backbone structures like VideoSwin [23], we do not shift the temporal windows across layers. To allow information propagation, we use very few blocks that can go across snippets. Inspired by ViT-Det [14], we evenly split a pre-trained ViT-based backbone into several subsets, and then insert a cross-snippet propagation module after the last block of each subset to enable information propagation. For the simplicity of the method, we adopt two simple but useful blocks as our cross-snippet propagation modules, coined as Local Propagation Block and Global Propagation Block.
@@
$\begin{split}q=xw_{q},k=xw_{k},v=xw_{v},\\ y=\mathrm{softmax}(qk^{T}/\sqrt{C_{q}})v\end{split}$ (1)
@@ 4.1 Ablation Study
In this section, we perform ablation experiments for ViT-TAD on THUMOS14 with the ViT-S backbone to discuss the architecture’s design in ViT-TAD.
@@
Table 1 ablates our inner-backbone information propagation strategy. These propagation blocks are placed in the backbone evenly and kernel size is set to $(3\times 3\times 3)$ for the local propagation block. We compare our global and local propagation blocks with baseline, both of them performs better than the baseline in 4-block settings. We further explore the number of propagation blocks and find that the local strategy benefits from having more blocks, but the global strategy does the opposite. There are two plausible explanations for this. First, the model’s ability to learn valuable information is hampered by frequent global interactions. Second, the self-attention operation cannot use prior information like the convolutional operator. To enhance temporal interaction, a more in-depth exploration is warranted.
@@
The design of global block is more flexible than local block and requires further discussion. As shown in Table 2a, additional clip-level positional encoding can enable the model to recognize clip-level feature sequences after snippet splicing, leading to better detection results. To save computational consumption, we disassemble the complete cross-snippet spatiotemporal relationship modeling into cross-snippet temporal modeling and inner-snippet spatiotemporal modeling. For vivid comparison, our global propagation block is referred to as 1D (temporal only) strategy shown in Fig 4(b), and applying clip-level spatiotemporal modeling on the last block of each subset in the backbone is referred to as 3D (temporal+spatial) strategy (Fig 4(a)). Both 1D and 3D strategies obtain similar detection results while 3D strategy causes more memory consumption shown in Table 2b, inferring that concentrating on clip-level temporal modeling is sufficient. We further study where global blocks should be located in the backbone. By default 4 global blocks are placed evenly. We compare by placing them in the first or last 4 blocks instead. As is shown in Table 2c, evenly 4 blocks configuration performs best, and last 4 blocks configuration follows. This is in line with the observation in ViT [10] that ViT has a longer attention distance in later blocks and is more localized in earlier ones. Premature temporal interaction of individual snippet features yields limited benefit. Finally, we adopt 4 evenly placed global blocks as the final configuration of the global propagation block.
@@
Table 3: Effectiveness of post-backbone information propagation. p-b is short for post-backbone information propagation.
prop. strategy use p-b 0.3 0.4 0.5 0.6 0.7 Avg
train infer
@@
Table 4: Study on different strategies using post-backbone information propagation. p-b is short for post-backbone information propagation. Here we explore how applying p-b or not during the training and inference phases affects the result.
@@
Post-backbone information propagation strategy is used for sufficient feature interaction among snippets. As shown in Table 3, only applying such strategy is also effective. Equipped with post-backbone information propagation, we then compare the detection results for both local and global blocks with the best configuration each in inner-backbone strategy. As is shown in Table 4, we could see that global propagation strategy obtains better detection results (from 62.3% to 63.0%), while the detection results of local propagation strategy drops (from 62.0% to 61.5%). This interesting finding prompts us to further investigate the impact on the modeling ability of post-backbone information propagation and backbone after introducing post-backbone information propagation. When we keep post-backbone information propagation during training but remove it during inference, detection results are worse compared with keeping it during inference for both global (from 63.0% to 62.2%) and local (from 61.5% to 59.4%) blocks. This confirms that post-backbone information propagation can improve the effect of global modeling. When both strategies discard it during inference and only compare the modeling capabilities of backbone, we find that backbone’s modeling ability of global propagation strategy is almost the same (compare 62.3% with 62.2%) but local propagation strategy drops a lot (from 62.0% to 59.4%). We propose that the varying modeling approaches employed by the convolutional operator and the self-attention mechanism contribute to this phenomenon. The former relies on priors to learn global features with translation invariance, while the latter disrupts such internal modeling, resulting in a diminished modeling capability of the backbone. Ultimately, we employ a global propagation strategy for both internal-backbone information propagation and post-backbone information propagation in the construction of our ViT-TAD.
Comparison between our ViT-TAD and other TAD pipelines.
@@
ActionFormer [43] ViT-S 75.2 70.0 62.0 48.7 33.4 57.9
ViT-S* 75.6 70.3 63.4 52.4 39.6 60.3
BasicTAD [42](baseline) ViT-S 74.6 70.2 62.6 51.3 38.4 59.4
ViT-TAD ViT-S 78.7 74.1 66.2 55.5 40.3 63.0
Table 5: Comparison between our ViT-TAD and other TAD pipelines. ViT-S* means that backbone feature is fine-tuned by BasicTAD in an end-to-end manner.
Method Backb. 0.3 0.4 0.5 0.6 0.7 Avg
TALLFormer Swin-B 76.0 - 63.2 - 34.5 59.2
@@
In this section, we compare ViT-TAD with three convincing methods of TAD, namely ActionFormer [43], BasicTAD [42] and TALLFormer [9]. ActionFormer is the state-of-the-art TAD pipeline that inputs pre-extracted temporal features for detection. BasicTAD is a simple end-to-end TAD pipeline whose head is adopted in ViT-TAD. Here we migrate the feature modeling capabilities of ViT-TAD to these methods to further explore the reasons why ViT-TAD achieves good detection results. Shown in Table 5, we use the feature extracted by ViT-S for ActionFormer, obtaining an average mAP of 57.9% on THUMOS14. While BasicTAD achieves an average mAP of 59.4%. To eliminate the benefit of end-to-end training manner, we treat the backbone fine-tuned by BasicTAD as a feature extractor for ActionFormer, making the detection accuracy of ActionFormer reach an average mAP of 60.3%. When we further adopt our proposed ViT-TAD, it can achieve an average mAP of 63.0%, even higher than ActionFormer with fine-tuned features from BasicTAD. Therefore we can draw several conclusions. First, the Actionformer’s detector is stronger than BasicTAD. Second, BasicTAD achieves better results after using the proposed modules in ViT-TAD, which further proves that ViT-TAD is a better TAD pipeline. In order to further verify the effectiveness of the module proposed in ViT-TAD, we migrate it to another end-to-end TAD method TAllFormer that handles clip-level temporal modeling. Shown in Table 6, our proposed modules improve TALLFormer’s detection results when both are applied with ViT-B (from 63.1% to 66.1%). This further proves the effectiveness of both inner-backbone and post-backbone information propagation strategies proposed in our ViT-TAD.
@@ 4.2 Analysis
In this section, we further analyze the efficiency and limitations of ViT-TAD.
@@
4 global blocks+p-b ViT-S112×${}_{112\times 112}$ 63.0 33.35M 9.33G 108.80G 2135
ViT-S160×${}_{160\times 160}$ 64.3 33.35M 23.56G(4.35G*) 220.59G 1349
ViT-B160×${}_{160\times 160}$ 69.0 131.25M 8.97G* 866.73G 845
Table 7: Practical performance of backbone adaptation strategies. p-b is short for post-backbone information propagation. *** means checkpoint training strategy is used.
We compare the differences between baseline and variants of ViT-TAD. As is shown in Table 7, ViT-TAD introduces additional training parameters and consumes more memory due to the introduction of multiple blocks in the backbone. Thanks to checkpoint training strategy (marked by *** in Table 7), it can still run successfully with limited computing resources.
@@
Figure 6: Heatmap visualization. (a): Original input video frames. (b): Heatmap from baseline method. (c): Heatmap from ViT-TAD.
@@ 4.4 Comparison with the State of the Art
PBRNet [19]96×${}_{96\times 96}$ I3D 58.5 54.6 51.3 41.8 29.5 47.1
AFSD [16]96×${}_{96\times 96}$ I3D 67.3 62.4 55.5 43.7 31.1 52.0
STPT [39]96×${}_{96\times 96}$ MViT 70.6 65.7 56.4 44.6 30.5 53.6
DaoTAD [35]112×${}_{112\times 112}$ res50-I3D 62.8 59.5 53.8 43.6 30.1 50.0
e2e-tadtr [20]112×${}_{112\times 112}$ SlowFast-R50 69.4 64.3 56.0 46.4 34.9 54.2
TALLFormer [9]224×${}_{224\times 224}$ Swin-B 76.0 - 63.2 - 34.5 59.2
TALLFormer [9]224×${}_{224\times 224}$ ViT-B 78.9 75.0 67.6 56.1 37.9 63.1
BasicTAD [42]160×${}_{160\times 160}$ SlowOnly-R50 75.5 70.8 63.5 50.9 37.4 59.6
[…]
ViT-TAD112×${}_{112\times 112}$ ViT-S 78.7 74.1 66.2 55.5 40.3 63.0
ViT-TAD160×${}_{160\times 160}$ ViT-S 79.8 75.2 68.4 56.4 41.7 64.3
ViT-TAD160×${}_{160\times 160}$ ViT-B 84.9 80.8 72.8 61.0 45.6 69.0
Table 8: Comparison with state-of-the-art methods on THUMOS14. The subscript indicates the spatial resolution for end-to-end methods. Flow denotes whether each method uses optical flow as input.
@@
STPT [39]96×${}_{96\times 96}$ MViT 51.40 33.70 6.80 33.40
AFSD [16]96×${}_{96\times 96}$ I3D 52.40 35.30 6.50 34.40
PBRNet [19]96×${}_{96\times 96}$ I3D 53.96 34.97 8.98 35.01
e2e-tadtr [20]112×${}_{112\times 112}$ SlowFast-R50 50.47 35.99 10.83 35.10
TALLFormer [9]224×${}_{224\times 224}$ Swin-B 54.10 36.20 7.90 35.60
[…]
ViT-TAD160×${}_{160\times 160}$ ViT-S 55.09 37.81 8.75 36.69
ViT-TAD160×${}_{160\times 160}$ ViT-B 55.44 37.91 8.99 37.12
Table 9: Comparison with state-of-the-art methods on ActivityNet-1.3. The subscript indicates the spatial resolution for end-to-end methods. Flow denotes whether each method uses optical flow as input.
@@
VideoMAE V2 [36] ViT-G - - - 18.20
BasicTAD [42]160×${}_{160\times 160}$ SlowOnly-R50 24.34 10.57 0.43 12.15
ViT-TAD160×${}_{160\times 160}$ ViT-B 32.61 15.85 2.68 17.20
Table 10: Comparison with state-of-the-art methods on FineAction. The subscript indicates the spatial resolution for end-to-end methods. Flow denotes whether each method uses optical flow as input.
We compare our ViT-TAD with the previous state-of-the-art methods on THUMOS14 [13], ActivityNet-1.3 [12] and FineAction [21]. We classify these methods into end-to-end methods (lower part of tables) and non-end-to-end methods (upper part of tables). For THUMOS14, the results are shown in Table 8. ViT-TAD with ViT-S backbone achieves mAP of 63.0% and obtains about 3.6% higher accuracy compared with baseline. We can also afford video frames with higher resolution and obtain average mAP of 64.3%. When we adopt ViT-B pre-trained from VideoMAE [32], our ViT-TAD can obtain an average mAP of 69.0% which outperforms the state-of-the-art end-to-end TAD methods by a large margin. For ActivityNet-1.3, the results are shown in Table 9. We attempt to adapt our model to predict binary action proposals and obtain the detection results by applying video-level action classifiers [40]. We get competitive results with ViT-S (36.69%) and state-of-the-art results with ViT-B (37.12%) among all end-to-end TAD methods. For FineAction, the results are shown in Table 10. we get good results with ViT-B (17.20%), only slightly lower than VideoMAE V2 [36] using ViT-G.

### v2

@@ Abstract
Vision Transformer (ViT) has shown high potential in video recognition, owing to its flexible design, adaptable self-attention mechanisms, and the efficacy of masked pre-training. Yet, it remains unclear how to adapt these pre-trained short-term ViTs for temporal action detection (TAD) in untrimmed videos. The existing works treat them as off-the-shelf feature extractors for each short-trimmed snippet without capturing the fine-grained relation among different snippets in a broader temporal context. To mitigate this issue, this paper focuses on designing a new mechanism for adapting these pre-trained ViT models as a unified long-form video transformer to fully unleash its modeling power in capturing inter-snippet relation, while still keeping low computation overhead and memory consumption for efficient TAD. To this end, we design effective cross-snippet propagation modules to gradually exchange short-term video information among different snippets from two levels. For inner-backbone information propagation, we introduce a cross-snippet propagation strategy to enable multi-snippet temporal feature interaction inside the backbone. For post-backbone information propagation, we propose temporal transformer layers for further clip-level modeling. With the plain ViT-B pre-trained with VideoMAE, our end-to-end temporal action detector (ViT-TAD) yields a very competitive performance to previous temporal action detectors, riching up to 69.5 average mAP on THUMOS14, 37.40 average mAP on ActivityNet-1.3 and 17.20 average mAP on FineAction.
✉: Corresponding author.
@@ 2 Related Work
Figure 2: Overview of ViT-TAD. We suppose the ViT-based backbone has $n$ blocks and divide them into several subsets. Each subset has $i$ blocks. We divide a video clip into several snippets and send each into the backbone for feature extraction. We perform temporal feature interaction among all snippets through the inner-backbone information propagation strategy. We further conduct clip-level modeling to refine clip-level features through the post-backbone information propagation strategy. * means the last layer is initialized as zero.
@@
Due to GPU memory constraints, most existing methods treated TAD [18, 19, 43, 47] as a two-step approach separating feature modeling and action detection. Since temporal features are extracted using backbones that are pre-trained on the action recognition dataset, there is a gap between these two different tasks, leading to sub-optimal performance. In order to address this gap, incorporating the backbone into the network’s training process forms an end-to-end TAD framework [17, 10, 20, 41, 46]. Simultaneously updating the network weights of both parts in the two-step approach will easily exceed the GPU memory limit. To address this issue, some methods [17, 20, 41] tried to downscale the frame resolution, while others tried to keep the resolution of the original video frame through improvements in training mechanisms. Among them, TALLFormer [10] attempted to build an offline memory bank to store features that cannot be updated synchronously and only update a portion of features during end-to-end training. Re2TAL [46] built a backbone with reversible modules to save memory consumption. The concurrent work AdaTAD [21] achieved the purpose of saving video memory by only training the adapter and freezing the backbone. Due to the existence of positional encoding, how to fine-tune the TAD model when spatiotemporal resolution differs from the pre-trained Transformer-based model is a problem. TALLFormer [10] and Re2TAL [46] circumvented this problem by keeping the same spatiotemporal resolution as the pre-trained backbone, while STPT [41] pre-trained its own backbone. We adopt the downscaling frame resolution strategy and use the operation from DeiT [35] to fix the positional encoding across resolutions to better use the pre-trained weight of the ViT-based backbone.
@@ 3.1 Inner-backbone Information Propagation
We design a cross-snippet propagation module for inner-backbone information exchange. Unlike hierarchical backbone structures like VideoSwin [25], we do not shift the temporal windows across layers. We use very few blocks that can go across snippets to allow information propagation. Inspired by ViT-Det [15], we evenly split a pre-trained ViT-based backbone into several subsets, and then insert a cross-snippet propagation module after the last block of each subset to enable information propagation. For the simplicity of the method, we adopt two simple but useful blocks as our cross-snippet propagation modules, coined as Local Propagation Block and Global Propagation Block.
@@
$\begin{split}q=xw_{q},k=xw_{k},v=xw_{v},\\ y=\mathrm{softmax}(qk^{T}/\sqrt{C_{q}})v,\end{split}$ (1)
@@ 4.1 Ablation Study
In this section, we perform ablation experiments for ViT-TAD on THUMOS14. Both ViT-S and ViT-B we adopted are pre-trained on Kinetics-400 [14] provided by [38].
@@
Table 1 ablates our inner-backbone information propagation strategy. These propagation blocks are placed in the backbone evenly and kernel size is set to $(3\times 3\times 3)$ for the local propagation block. We compare our global and local propagation blocks with the baseline, and both of them perform better than the baseline in 4-block settings. We further explore the number of propagation blocks and find that the local strategy benefits from having more blocks, but the global strategy does the opposite. There are two plausible explanations for this. First, the model’s ability to learn valuable information is hampered by frequent global interactions. Second, the self-attention operation cannot use prior information like the convolutional operator. To enhance temporal interaction, a more in-depth exploration is warranted.
@@
The design of the global block is more flexible than the local block and requires further discussion. As shown in Table 2a, additional clip-level positional encoding can enable the model to recognize clip-level feature sequences after snippet splicing, leading to better detection results. To save computational consumption, we disassemble the complete cross-snippet spatiotemporal relationship modeling into cross-snippet temporal modeling and inner-snippet spatiotemporal modeling. For vivid comparison, our global propagation block is referred to as 1D (temporal only) strategy shown in Fig 4(b), and applying clip-level spatiotemporal modeling on the last block of each subset in the backbone is referred to as 3D (temporal+spatial) strategy (Fig 4(a)). Both 1D and 3D strategies obtain similar detection results while 3D strategy causes more memory consumption shown in Table 2b, inferring that concentrating on clip-level temporal modeling is sufficient. We further study where global blocks should be located in the backbone. By default 4 global blocks are placed evenly. We compare by placing them in the first or last 4 blocks instead. As is shown in Table 2c, evenly 4 blocks configuration performs best, and last 4 blocks configuration follows. This is in line with the observation in ViT [11] that ViT has a longer attention distance in later blocks and is more localized in earlier ones. Premature temporal interaction of individual snippet features yields limited benefit. Finally, we adopt 4 evenly placed global blocks as the final configuration of the global propagation block.
@@
Table 3: Study on Post-backbone Information Propagation. p-b is short for post-backbone information propagation.
@@
Post-backbone information propagation strategy is used for sufficient feature interaction among snippets. It shows that only applying such a strategy (”none” in Table 3) is also effective. Equipped with post-backbone information propagation, we then compare the detection results for both local and global blocks with the best configuration each in the inner-backbone strategy. As is shown in Table 3, we could see that the global propagation strategy obtains better detection results (from 62.3% to 63.0%), while the detection results of the local propagation strategy drop (from 62.0% to 61.5%), so we employ the global propagation strategy for both inner-backbone and post-backbone information propagation strategies.
Comparison between ViT-TAD and other TAD Pipelines.
Method details 0.3 0.4 0.5 0.6 0.7 Avg
[…]
ViT-S* 75.6 70.3 63.4 52.4 39.6 60.3
[…]
Table 4: Comparison between ViT-TAD and other TAD pipelines. ViT-S* means BasicTAD fine-tunes that backbone feature in an end-to-end manner. Details include the choice of backbone, frame resolution, and frame rate.
@@
TALLFormer [10] Swin-B 76.0 - 63.2 - 34.5 59.2
@@
In this section, we compare ViT-TAD with four convincing methods of TAD, namely ActionFormer [45], BasicTAD [44], TALLFormer [10] and Tridet [31]. Here we migrate the feature modeling capabilities of ViT-TAD to these methods to further explore the reasons why ViT-TAD achieves good detection results. Shown in Table 4, we use the feature extracted by ViT-S for ActionFormer, obtaining an average mAP of 57.9% on THUMOS14. While BasicTAD achieves an average mAP of 59.4% with a lower frame rate and resolution. To eliminate the benefit of end-to-end training manner, we treat the backbone fine-tuned by BasicTAD as a feature extractor for ActionFormer, making the detection accuracy of ActionFormer reach an average mAP of 60.3%. When we further adopt our ViT-TAD based on BasicTAD, it can achieve an average mAP of 63.0%, even higher than ActionFormer with fine-tuned features from BasicTAD. When we further compare our ViT-TAD with the state-of-the-art TAD method TriDet that inputs pre-extracted features, its performance is worse than our ViT-TAD when using pre-extracted ViT-B features based on 30 FPS frame rate and $224\times 224$ frame resolution (compare 68.2% with 69.5%), and even worse under fair input conditions (compare 60.0% with 69.5%). Therefore we can draw several conclusions. First, the Actionformer’s detector is stronger than BasicTAD. Second, BasicTAD achieves better results after using the proposed modules in ViT-TAD. Third, ViT-TAD outperforms the state-of-the-art TAD method TriDet under fair input conditions. To further verify the module’s effectiveness proposed in ViT-TAD, we migrate it to another end-to-end TAD method TALLFormer that handles clip-level temporal modeling. Shown in Table 5, our proposed modules improve TALLFormer’s detection results when both are applied with ViT-B provided by [38] (from 63.1% to 66.1%). In summary, ViT-TAD has significant advantages over the above TAD methods, and its performance is better than the current non-end-to-end TAD method TriDet. At the same time, the modules proposed in ViT-TAD can be adopted by other TAD methods to achieve better detection results.
Runtime Comparison with other TAD Methods.
Methods details GPU FPS Avg
[…]
Table 6: Runtime comparison with other TAD methods. Details include the choice of backbone, frame resolution, and frame rate.
Here we list the runtime comparison with other TAD methods shown in Table 6. The time cost of optical flow extraction makes PBRNet [20] and AFSD [17] lower than the reported speed. TriDet [31] and ActionFormer [45] need to extract features of the complete video in advance, so we consider the inference speed of both feature extraction and model inference. Since our method is based on ViT, its inference speed cannot be compared with convolution-based methods [22, 44], but our ViT-TAD achieves the best detection results while ensuring fast inference speed among methods using ViT.
@@ 4.2 Analysis
[…]
Table 7: Practical performance of backbone adaptation strategies. p-b is short for post-backbone information propagation. Details include backbone, frame resolution, and frame rate. * means checkpoint training strategy is used.
We compare the differences between baseline and variants of ViT-TAD. As is shown in Table 7, ViT-TAD introduces additional training parameters and consumes more memory due to the introduction of multiple blocks in the backbone. Thanks to checkpoint training strategy (marked by * in Table 7), it can still run successfully with limited computing resources.
@@
Figure 6: Heatmap visualization. (a): Original input frames. (b): Heatmap from baseline method. (c): Heatmap from ViT-TAD.
@@ 4.4 Comparison with the State of the Art
VideoMAE V2 [38] ViT-G 84.0 79.6 73.0 63.5 47.7 69.6
PBRNet [20]96×96 I3D 58.5 54.6 51.3 41.8 29.5 47.1
AFSD [17]96×96 I3D 67.3 62.4 55.5 43.7 31.1 52.0
STPT [41]96×96 MViT 70.6 65.7 56.4 44.6 30.5 53.6
DaoTAD [37]112×112 res50-I3D 62.8 59.5 53.8 43.6 30.1 50.0
e2e-tadtr [22]112×112 SlowFast-R50 69.4 64.3 56.0 46.4 34.9 54.2
TALLFormer [10]224×224 Swin-B 76.0 - 63.2 - 34.5 59.2
TALLFormer [10]224×224 ViT-B 78.9 75.0 67.6 56.1 37.9 63.1
BasicTAD [44]160×160 SlowOnly-R50 75.5 70.8 63.5 50.9 37.4 59.6
Re2TAL [46]224×224 Swin-T 77.0 71.5 62.4 49.7 36.3 59.4
Re2TAL [46]224×224 SlowFast-R101 77.4 72.6 64.9 53.7 39.0 61.5
BasicTAD [44](baseline)112×112 ViT-S 74.6 70.2 62.6 51.3 38.4 59.4
ViT-TAD112×112 ViT-S 78.7 74.1 66.2 55.5 40.3 63.0
ViT-TAD160×160 ViT-S 79.8 75.2 68.4 56.4 41.7 64.3
ViT-TAD160×160 ViT-B 85.1 80.9 74.2 61.8 45.4 69.5
Table 8: Comparison with state-of-the-art methods on THUMOS14. The subscript indicates the spatial resolution. Flow denotes whether each method uses optical flow as input.
@@
STPT [41]96×96 MViT 51.40 33.70 6.80 33.40
AFSD [17]96×96 I3D 52.40 35.30 6.50 34.40
PBRNet [20]96×96 I3D 53.96 34.97 8.98 35.01
e2e-tadtr [22]112×112 SlowFast-R50 50.47 35.99 10.83 35.10
TALLFormer [10]224×224 Swin-B 54.10 36.20 7.90 35.60
Re2TAL [46]224×224 Swin-T 54.75 37.81 9.03 36.80
Re2TAL [46]224×224 SlowFast-R101 55.25 37.86 9.05 37.01
ViT-TAD160×160 ViT-S 55.09 37.81 8.75 36.69
ViT-TAD160×160 ViT-B 55.87 38.47 8.80 37.40
Table 9: Comparison with state-of-the-art methods on ActivityNet-1.3. The subscript indicates the spatial resolution. Flow denotes whether each method uses optical flow as input.
@@
VideoMAE V2 [38] ViT-G 29.07 17.66 5.07 18.24
BasicTAD [44]160×160 SlowOnly-R50 24.34 10.57 0.43 12.15
ViT-TAD160×160 ViT-B 32.61 15.85 2.68 17.20
Table 10: Comparison with state-of-the-art methods on FineAction. The subscript indicates the spatial resolution. Flow denotes whether each method uses optical flow as input.
We compare our ViT-TAD with the previous state-of-the-art methods on THUMOS14 [13], ActivityNet-1.3 [6] and FineAction [23]. We classify these methods into end-to-end methods (lower part of tables) and non-end-to-end methods (upper part of tables). For THUMOS14, the results are shown in Table 8. Our ViT-TAD with ViT-B outperforms all TAD methods except VideoMAE V2 which adopts ViT-G. For ActivityNet-1.3, the results are shown in Table 9. We attempt to adapt our model to predict binary action proposals and obtain the detection results by applying video-level action classifiers [42]. We get competitive results with ViT-S (36.69%) and state-of-the-art results with ViT-B (37.40%) among all end-to-end TAD methods. For FineAction, the results are shown in Table 10. We get good results with ViT-B (17.20%), only slightly lower than VideoMAE V2 [38].